OpenStudy (savannah_noelle):

I'm trying to figure out what the answer would be to this convex equation, can someone help me please? The convex lens has a focal length of 10.0cm, and the object distance is 15.0cm. The formula is 1/d(o) + 1/d(i) = 1/f. I got 0.06666667, is that correct?

OpenStudy (anonymous):

\[\frac{ 1 }{ i }+\frac{ 1 }{ 15 }=\frac{ 1 }{ 10 },\] \[\frac{ 1 }{ i }=\frac{ 1 }{ 10 }-\frac{ 1 }{ 15 }=\frac{ 2-1 }{ 30 }=\frac{ 1 }{ 30 }\] distance of image from the lens=30 cm

OpenStudy (savannah_noelle):

Thank you! Okay so if I replaced the number "15" with "20" would the answer be 20cm or is that wrong?

OpenStudy (anonymous):

that is correct.

OpenStudy (savannah_noelle):

Yayy, thank you! I just have one more question. How do you know the characteristics of a convex and concave lens? Also, do concave and convex lenses use the same equations?

OpenStudy (savannah_noelle):

I'm filling out this chart on convex and concave lenses. I can do the math part now (thanks to you!) but I still can't figure out how to know if it's small,large, inverted, upright, real, or virtual. This is what my chart looks like. @surjithayer

OpenStudy (anonymous):

simple thing for convex lens to remember. i denote pole =o focal length =f distance of object from pole=u distance of image from pole=v if the object is 1. at f,image at infinity. 2. between f and 2 f image is away from 2 f. 3. at 2 f ,image at 2 f. 4.away 2f ,image between f and 2 f. 5. at infinity image at f all these images are inverted and real remember for a convex lens distance between object and image is \[\ge 4f \]

OpenStudy (savannah_noelle):

I'm still confused.

OpenStudy (savannah_noelle):

so for the first equation we did would it be smaller, upright, and real? @surjithayer